# Asymptotic condition numbers for linear ODEs

### v1

@@ 2. Linear Algebra Results
We introduce linear algebra notations and results necessary for the analysis developed in this paper. Although the content pertains to the well-known topic of the JCF of a matrix, it addresses very specific aspects that are either not known or not sufficiently detailed for the purposes of this paper.
This section can be skipped and referred back to as needed while reading the subsequent sections.
@@ 2.1. The JCF of the matrix $A$
  • For any $i\in\{1,\ldots,p\}$ and for any $j\in\{1,\ldots,d_{i}\}$, the Jordan mini-block $J^{\left(i,j\right)}$ is upper bi-diagonal:
@@
An eigenvalue $\lambda_{i}$, $i\in\{1,\ldots,p\}$, is called defective if $d_{i}<\nu_{i}$ and non-defective if $d_{i}=\nu_{i}$. Clearly, $\lambda_{i}$ is non-defective if and only if $m_{i}=1$, i.e. $J^{(i)}$ is diagonal. The matrix $A$ is diagonalizable if and only if all the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ are non-defective, i.e. $J$ is diagonal.
@@
2.1.1. Three-index notation
@@
  • $i\in\{1,\ldots,p\}$ is the index of the block $J^{(i)}$ traversed by the $l$-th column (or row) of $J$.;
@@
  • $k\in\left\{1,\ldots,m_{ij}\right\}$ is the index of the column (or row) of the mini-block $J^{(i,j)}$ included in the $l$-th column (row) of $J$.
@@ 2.1.2. The Jordan basis
2.1.3. Jordan chains
The Jordan basis (8) is partitioned in the Jordan chains
@@ 2.1.3. Jordan chains
has not a solution.
Observe that (7) is equivalent to have (2.1.3), for $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, and that there is a correspondence one-to-one between mini-blocks and Jordan chains.
2.1.4. The matrix $zA$
@@
Let $z\in\mathbb{C}\setminus\{0\}$. The distinct eigenvalues of $zA$ are $z\lambda_{i}$$i\in\{1,\ldots,p\}$. For any $i\in\{1,\ldots,p\}$, the number and the dimensions of the mini-blocks corresponding to the eigenvalue $z\lambda_{i}$ in the JCF of $zA$ are equal to the number and the dimensions of the mini-blocks corresponding to the eigenvalue $\lambda_{i}$ in the JCF of $A$. A Jordan basis of $zA$ is
@@
By exchanging the role of $A$ and $zA$, i.e. we consider $zA$ and $z^{-1}(zA)$, we also see that given a Jordan chain
@@
is a Jordan chain of $A$ corresponding to the eigenvalue $\lambda_{i}$ (here, $m_{ij}(zA)$ denotes the length of a Jordan chain of $zA$). This shows that the Jordan chains of $zA$ are all of type (11). Consequently, the number and the lenghts of the Jordan chains corresponding to the eigenvalue $z\lambda_{i}$ of $\alpha A$ are equal to the number and the lengths of the Jordan chains corresponding to the eigenvalue $\lambda_{i}$ of $A$. Therefore, the number and the dimensions of the mini-blocks corresponding to the eigenvalue $z\lambda_{i}$ of $zA$ are equal to the number and the dimensions of the mini-blocks corresponding to the eigenvalue $\lambda_{i}$ of $A$.
@@ 2.1.5. The real case
In fact, by conjugating both sides in (2.1.3), we have
@@
By exchanging the role of $\lambda_{i_{1}}$ and $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, i.e. we consider $\lambda_{i_{2}}$ and $\overline{\lambda_{i_{2}}}$, we also see that given a Jordan chain
@@
is a Jordan chain corresponding to the eigenvalue $\lambda_{i_{1}}=\overline{\lambda_{i_{2}}}$. This shows that the Jordan chains corresponding to $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ are all of type (14). This implies (12). By collecting all the Jordan chains (14) corresponding to $\lambda_{i_{1}}$, we obtain a Jordan basis satisfying (13).
@@
  • the eigenspace of $\lambda_{i}$ in $\mathbb{C}^{n}$ has a basis of real eigenvalues;
@@
2.2. The matrices $V^{(i)}$, $W^{(i)}$, $N^{(i,l)}$ and the vector $\alpha^{(i)}(u)$
@@
Remind that $V$ is the matrix whose columns (8) constitute a Jordan basis.
@@
    be the rows of $W$ in the three-index notation. The $n$ row vectors in (16) are called left generalized eigenvectors of $A$, whereas, as we have already seen, the $n$ columns vectors in (8) are called (right) generalized eigenvectors of $A$. For $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, let
@@
    $\alpha^{(i,j)}(u):=\left[\begin{array}[]{c}\alpha_{ij1}(u)\\ \vdots\\ \alpha_{ijm_{ij}}(u)\end{array}\right]\in\mathbb{C}^{m_{ij}}.$
@@
    Observe that the matrix $N^{(i,l)}$ has the same dimensions $\nu_{i}\times\nu_{i}$ of the Jordan block $J^{(i)}=\mathrm{diag}\left(J^{(i,1)},\ldots,J^{(i,d_{i})}\right)$ and the diagonal blocks $N^{(i,l,1)},$ $\ldots$,… , $N^{(i,l,d_{i})}$ have the same dimensions of the Jordan mini-blocks $J^{(i,1)},\ldots,$ $J^{(i,d_{i})}$, respectively. Thus, as it is illustrated in Figure 2, the matrix $N^{(i,l)}$ has $1$ in the intersection of the $l$-th upper diagonal with the (frame of the) Jordan mini-blocks and 00 in all the other places.
@@
2.2.1. The matrix $zA$
When we replace $A$ by $zA$, $z\in\mathbb{C}\setminus\{0\}$, the number and the dimensions of blocks and mini-blocks remain the same (remind Proposition 2). Moreover, a Jordan basis for $zA$ is given in (10). Therefore, we know how it is transformed the matrix $V$ by replacing $A$ by $zA$. The next proposition says this and, in addition, how it is transformed the matrix $W=V^{-1}$.
@@ 2.2.2. The real case
Assume $A\in\mathbb{R}^{n\times n}$ and $u\in\mathbb{R}^{n}$. Moreover, assume to have a Jordan basis as that described in Proposition 3.
@@
Since we are assuming to have a Jordan basis as that described in Proposition 3, we have $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$.
Now, we prove that for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ we have $\alpha^{(i)}(u)\in\mathbb{R}$, and for $i_{1},i_{2}\in\{1,\ldots,p\}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ with $\mathrm{Im}\left(\lambda_{i}\right)>0$ we have $\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}$. By conjugating both sides of
@@
So, by separating real eigenvalues and complex conjugate pairs of eigenvalues, we have
@@
by reminding that $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $\lambda_{i_{1}}\in\mathbb{C}$ with $\mathrm{Im}\left(\lambda_{i}\right)>0$. Since $u$ can be expressed as a linear combination of the Jordan basis in a unique manner, we obtain $\alpha^{(i)}(u)=\overline{\alpha^{(i)}(u)}$, i.e. $\alpha^{(i)}(u)\in\mathbb{R}$, for $\lambda_{i}\in\mathbb{R}$ and $\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}$ for $\lambda_{i_{1}}\in\mathbb{C}$ with $\mathrm{Im}\left(\lambda_{i_{1}}\right)>0$.
@@
2.3. The formula for $\mathrm{e}^{tA}$
Next proposition provides the announced formula for the matrix exponential $\mathrm{e}^{tA}$. For taste of generality, we consider the general situation of a function of matrix $f(A)$, where $f(z)$, $z\in\mathcal{D}\subseteq\mathbb{C}$, is an analytic complex function of $z$. The domain $\mathcal{D}$ of $f$ is an open subset of $\mathbb{C}$ and we assume that the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ are contained in $\mathcal{D}$.
@@ 2.3. The formula for $\mathrm{e}^{tA}$
By recalling the definition of function of matrix by the JCF (see, e.g., [8]), we have
@@
2.3.1. The matrices $P_{i0}$
For $i\in\{1,\ldots,p\}$, since $N^{(i,0)}=I_{n}$, we have
@@ 2.3.1. The matrices $P_{i0}$
The matrix $P_{i0}$ is the projection onto the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e. the subspace spanned by the generalized eigenvectors corresponding to the eigenvalue $\lambda_{i}$. In fact, we have, for $u\in\mathbb{C}^{n}$,
@@ 2.3.2. The case $A$ diagonalizable
In this case of $A$ diagonalizable, the matrix $P_{i0}$ is the projection onto the eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e. the subspace spanned by the eigenvectors corresponding to the eigenvalue $\lambda_{i}$.
@@ 2.4.1. An alternative expression for $P_{il}$
in (29) and (30) are the indeces $j$ of the mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal. The indeces $k$ in the inner sum
@@
are the row indeces of the elements of the mini-block $J^{(i,j)}$ on this $l$-th upper diagonal (see Figure 3). The column indeces $l+k$ of these elements appear as third indeces in $w^{(i,j,l+k)}$ in (29) and $\alpha_{ij\,l+k}(u)$ in (30). In other words, with reference to the previous Figure 2, in the double sum in (29) and (30) we are summing over all elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with the mini-blocks: the three-index notation $(i,j,k)$ of the row of these elements appears in $v^{(i,j,k)}$ and the three-index notation $(i,j,l+k)$ of the columns appears in $w^{(i,j,l+k)}$ for (29) and in $\alpha_{ijl+k}(u)$ for (30).
@@
As for (29), by the definition (23) we have
$P_{il}$ $=$ $\sum\limits_{j=1}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}$
$=$ $\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}$
@@
by reminding the form (18) of $N^{(i,l,j)}$.
As for (30), we have
@@
Figure 3. Indeces $j$ and $k$ in (29) and (30): $j$ is an index of mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal and $k$ is an index of row of elements in the $l$-th upper diagonal.
@@
    If there exists an unique $j\in\{1,\ldots,d_{i}\}$ such that $m_{i}=m_{ij}$, then
@@
    since these vectors are multiplied by coefficients different from the corresponding components $\alpha_{ijk}(u)$. The coefficients are the components $\alpha_{ijl+k}(u)$ shifted by $l$ in the third index.
2.4.2. The matrix $zA$
The next proposition shows how the matrices $P_{il}$ are transformed when the matrix $A$ is replaced by $zA$, $z\in\mathbb{C}\setminus\{0\}$. Observe that, when $A$ is replaced by $zA$, the number and the dimensions of blocks and mini-blocks remain the same (remind Proposition 2). Thus, the indeces $i$ and $l$ for the matrices $P_{il}(zA)$ corresponding to $zA$ are the same indeces $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$ for the matrices $P_{il}$ corresponding to $A$.

### v2

@@ Appendix A Linear Algebra Results
We introduce linear algebra notations and results necessary for the analysis developed in this paper. Although the content pertains to the well-known topic of the Jordan Canonical Form (JCF) of a matrix, it addresses very specific aspects that are either not known or not sufficiently detailed for the purposes of this paper.
@@ A.1. The JCF of the matrix AA
  • For any $i\in\{1,\ldots,p\}$ and for any $j\in\{1,\ldots,d_{i}\}$, the Jordan mini-block $J^{\left(i,j\right)}$ is upper bidiagonal:
@@
An eigenvalue $\lambda_{i}$, $i\in\{1,\ldots,p\}$, is called defective if $d_{i}<\nu_{i}$ and non-defective if $d_{i}=\nu_{i}$. Clearly, $\lambda_{i}$ is non-defective if and only if $m_{i}=1$, i.e. $J^{(i)}$ is diagonal. The matrix $A$ is diagonalizable if and only if all the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ are non-defective, i.e., $J$ is diagonal.
@@
A.1.1. Three-index notation
@@
  • $i\in\{1,\ldots,p\}$ is the index of the block $J^{(i)}$ traversed by the $l$-th column (or row) of $J$;
@@
  • $k\in\left\{1,\ldots,m_{ij}\right\}$ is the index of the column (or row) of the mini-block $J^{(i,j)}$ included in the $l$-th column (or row) of $J$.
@@ A.1.2. The Jordan basis
A.1.3. Jordan chains
The Jordan basis (31) is partitioned into the Jordan chains
@@ A.1.3. Jordan chains
has no solution.
Observe that (30) is equivalent to have (A.1.3), for all $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, and that there is a correspondence one-to-one between mini-blocks and Jordan chains.
A.1.4. The matrix zAzA
@@
Let $z\in\mathbb{C}\setminus\{0\}$. The distinct eigenvalues of $zA$ are $z\lambda_{i}$, $i\in\{1,\ldots,p\}$. For any $i\in\{1,\ldots,p\}$, the number and the dimensions of the mini-blocks corresponding to the eigenvalue $z\lambda_{i}$ in the JCF of $zA$ are equal to the number and the dimensions of the mini-blocks corresponding to the eigenvalue $\lambda_{i}$ in the JCF of $A$. A Jordan basis of $zA$ is
@@
By exchanging the role of $A$ and $zA$, i.e., we consider $zA$ and $z^{-1}(zA)$, we also see that, given a Jordan chain
@@
is a Jordan chain of $A$ corresponding to the eigenvalue $\lambda_{i}$. Here, $m_{ij}(zA)$ denotes the length of a Jordan chain of $zA$. This shows that the Jordan chains of $zA$ are all of type (34). Consequently, the number and the lenghts of the Jordan chains corresponding to the eigenvalue $z\lambda_{i}$ of $zA$ are equal to the number and the lengths of the Jordan chains corresponding to the eigenvalue $\lambda_{i}$ of $A$. Therefore, the number and the dimensions of the mini-blocks corresponding to the eigenvalue $z\lambda_{i}$ of $zA$ are equal to the number and the dimensions of the mini-blocks corresponding to the eigenvalue $\lambda_{i}$ of $A$.
@@ A.1.5. The real case
In fact, by conjugating both sides in (A.1.3), we have
@@
By exchanging the role of $\lambda_{i_{1}}$ and $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$, i.e., we consider $\lambda_{i_{2}}$ and $\overline{\lambda_{i_{2}}}$, we also see that, given a Jordan chain
@@
is a Jordan chain corresponding to the eigenvalue $\lambda_{i_{1}}=\overline{\lambda_{i_{2}}}$. This shows that the Jordan chains corresponding to $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ are all of type (37). This implies (35). By collecting all the Jordan chains (37) corresponding to $\lambda_{i_{2}}$, we obtain a Jordan basis satisfying (36).
@@
  • the eigenspace of $\lambda_{i}$ in $\mathbb{C}^{n}$ has a basis of real eigenvectors;
@@
A.2. The matrices $V^{(i)}$, $W^{(i)}$, $N^{(i,l)}$ and the vector $\alpha^{(i)}(u)$
@@
Recall that $V$ is the matrix whose columns (31) constitute a Jordan basis.
@@
    be the rows of $W$ in the three-index notation. They appear in $W$ in lexicographic order. The $n$ row vectors in (39) are called left generalized eigenvectors of $A$, whereas, as we have already seen, the $n$ column vectors in (31) are called (right) generalized eigenvectors of $A$. For $i\in\{1,\ldots,p\}$ and $j\in\{1,\ldots,d_{i}\}$, let
@@
    $\alpha^{(i,j)}(u):=\left[\begin{array}[]{c}\alpha_{ij1}(u)\\ \vdots\\ \alpha_{ijm_{ij}}(u)\end{array}\right]\in\mathbb{C}^{m_{ij}}$
@@
    Observe that the matrix $N^{(i,l)}$ has the same dimensions $\nu_{i}\times\nu_{i}$ of the Jordan block $J^{(i)}=\mathrm{diag}\left(J^{(i,1)},\ldots,J^{(i,d_{i})}\right)$ and the diagonal blocks $N^{(i,l,1)},$ $\ldots,$ $N^{(i,l,d_{i})}$ have the same dimensions of the Jordan mini-blocks $J^{(i,1)},\ldots,$ $J^{(i,d_{i})}$, respectively. Thus, as it is illustrated in Figure 6, the matrix $N^{(i,l)}$ has $1$ in the intersection of the $l$-th upper diagonal with the (frame of the) Jordan mini-blocks and 0 in all the other places.
Remark 23.
Regarding the matrices $N^{(i,l,j)}$ and $N^{(i,l)}$, where $i\in\{1,\dots,p\}$, $l\in\{1,\ldots,m_{i}-1\}$ and $j\in\{1,\ldots,d_{i}\}$, note that
$N^{(i,l,j)}=\left(N^{(i,1,j)}\right)^{l}\text{\ \ and\ \ }N^{(i,l)}=\left(N^{(i,1)}\right)^{l}.$
Viewing $N^{(i,l,j)}$ and $N^{(i,l)}$ as powers of the nilpotent matrices $N^{(i,1,j)}$ and $N^{(i,1)}$ is how the matrices $N^{(i,l,j)}$ and $N^{(i,l)}$ are presented in usual expositions of the JCF.
@@
A.2.1. The matrix zAzA
When we replace $A$ by $zA$, $z\in\mathbb{C}\setminus\{0\}$, the number and the dimensions of blocks and mini-blocks remain the same (recall Proposition 21). Moreover, a Jordan basis for $zA$ is given in (33). Therefore, we know how the matrix $V$ is transformed by replacing $A$ by $zA$. The next proposition says this and, in addition, how the matrix $W=V^{-1}$ is transformed.
@@ A.2.2. The real case
Assume $A\in\mathbb{R}^{n\times n}$ and $u\in\mathbb{R}^{n}$. Moreover, assume we have a Jordan basis as that described in Proposition 22.
@@
Since we have a Jordan basis as that described in Proposition 22, we have $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $i_{1},i_{2}\in\{1,\ldots,p\}$ with $i_{1}\neq i_{2}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$.
Now, we prove that for $i\in\{1,\ldots,p\}$ such that $\lambda_{i}\in\mathbb{R}$ we have $\alpha^{(i)}(u)\in\mathbb{R}$, and for $i_{1},i_{2}\in\{1,\ldots,p\}$ such that $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ with $\mathrm{Im}\left(\lambda_{i_{1}}\right)>0$ we have $\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}$. By conjugating both sides of
@@
Therefore, by separating real eigenvalues and complex conjugate pairs of eigenvalues, we have
@@
by recalling that $V^{(i)}\in\mathbb{R}^{n\times\nu_{i}}$ for $\lambda_{i}\in\mathbb{R}$ and $V^{(i_{2})}=\overline{V^{(i_{1})}}$ for $\lambda_{i_{1}}\in\mathbb{C}$ with $\mathrm{Im}\left(\lambda_{i}\right)>0$. Since $u$ can be expressed as a linear combination of the Jordan basis in a unique manner, we obtain $\alpha^{(i)}(u)=\overline{\alpha^{(i)}(u)}$, i.e. $\alpha^{(i)}(u)\in\mathbb{R}$, for $\lambda_{i}\in\mathbb{R}$ and $\alpha^{(i_{2})}(u)=\overline{\alpha^{(i_{1})}(u)}$ for $\lambda_{i_{1}}\in\mathbb{C}$ with $\mathrm{Im}\left(\lambda_{i_{1}}\right)>0$.
@@
A.3. The formula for $\mathrm{e}^{tA}$
Next proposition provides the announced formula for the matrix exponential $\mathrm{e}^{tA}$. For sake of generality, we consider a matrix function $f(A)$, where $f(z)$, $z\in\mathcal{D}\subseteq\mathbb{C}$, is an analytic complex function of $z$. The domain $\mathcal{D}$ of $f$ is an open subset of $\mathbb{C}$ and we assume that the eigenvalues $\lambda_{1},\ldots,\lambda_{p}$ of $A$ are contained in $\mathcal{D}$.
@@ A.3. The formula for $\mathrm{e}^{tA}$
By recalling the definition of matrix function by the JCF (see, e.g., [7]), we have
@@
A.3.1. The matrices $P_{i0}$
For $i\in\{1,\ldots,p\}$, since $N^{(i,0)}=I_{\nu_{i}}$, we have
@@ A.3.1. The matrices $P_{i0}$
The matrix $P_{i0}$ is the projection onto the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e., the subspace spanned by the generalized eigenvectors corresponding to the eigenvalue $\lambda_{i}$. In fact, we have, for $u\in\mathbb{C}^{n}$,
@@ A.3.2. The case $A$ diagonalizable
In this case of $A$ diagonalizable, the matrix $P_{i0}$ is the projection onto the eigenspace corresponding to the eigenvalue $\lambda_{i}$, i.e., the subspace spanned by the eigenvectors corresponding to the eigenvalue $\lambda_{i}$.
@@ A.4.1. An alternative expression for $P_{il}$
in (52) and (53) are the indices $j$ of the mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal. The indices $k$ in the inner sum
@@
are the row indices of the elements of the mini-block $J^{(i,j)}$ on this $l$-th upper diagonal (see Figure 7). The column indices $l+k$ of these elements appear as third indices in $w^{(i,j,l+k)}$ in (52) and $\alpha_{ij\,l+k}(u)$ in (53).
In other words, with reference to the previous Figure 6, in the double sum in (52) and (53) we are summing over all elements in the intersection of the $l$-th diagonal of $N^{(i,l)}$ with the mini-blocks: the three-index notation $(i,j,k)$ of the row index of these elements appears in $v^{(i,j,k)}$ and the three-index notation $(i,j,l+k)$ of the column index appears in $w^{(i,j,l+k)}$ for (52) and in $\alpha_{ijl+k}(u)$ for (53).
@@
Regarding (52), by the definition (46) we have
$P_{il}$ $=$ $\sum\limits_{j=1}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}=\sum\limits_{\begin{subarray}{c}j=1\\ m_{ij}\geq l+1\end{subarray}}^{d_{i}}V^{\left(i,j\right)}N^{(i,l,j)}W^{\left(i,j\right)}$
@@
by recalling the form (41) of $N^{(i,l,j)}$.
Regarding (53), we have
@@
Figure 7. Indices $j$ and $k$ in (52) and (53): $j$ is an index of mini-blocks $J^{(i,j)}$ of $J^{(i)}$ having the $l$-th upper diagonal and $k$ is an index of row of elements in the $l$-th upper diagonal.
@@
    If there exists a unique $j\in\{1,\ldots,d_{i}\}$ such that $m_{i}=m_{ij}$, then
@@
    since these vectors are multiplied by coefficients different from the corresponding components $\alpha_{ijk}(u)$. These coefficients are the components $\alpha_{ijl+k}(u)$ shifted by $l$ in the third index.
A.4.2. The matrix zAzA
The next proposition shows how the matrices $P_{il}$ are transformed when the matrix $A$ is replaced by $zA$, $z\in\mathbb{C}\setminus\{0\}$. Observe that, when $A$ is replaced by $zA$, the number and the dimensions of blocks and mini-blocks remain the same (see Proposition 21). Thus, the indices $i$ and $l$ for the matrices $P_{il}(zA)$ corresponding to $zA$ are the same indices $i\in\{1,\ldots,p\}$ and $l\in\{0,\ldots,m_{i}-1\}$ for the matrices $P_{il}$ corresponding to $A$.